OpenStudy (anonymous):

If two angles are supplementary (sum of the two angles equals 180 degrees), find each angle is one angle is 30 degrees more than twice the other angle.

OpenStudy (anonymous):

Let one angle be x. Then, other angle is 2x+30. x+2x+30=180 3x+30=180 3x=150 x=50 degrees. 2x+30=50*2+30 =130 degrees.